# Ground state properties of the Heisenberg-compass model on the square lattice

### v1

@@ VI Quantum ground state phase diagram
We have described above the classical phase diagram of the Heisenberg-compass model using several complementary methods: Luttinger-Tisza analysis, low-temperature spin-wave expansion and Monte Carlo simulations. In this section, we explore the quantum ground state phase diagram of this model. To begin, we examine the role of quantum fluctuations on the classical ground states at zero temperature. We commit a particular attention in the regions where the classical model displays an accidental degeneracy and quantum ObD is expected to determine the ordering pattern.
@@ VI.1 Order by disorder from quantum fluctuations at $T=0$
We start by considering the effects of quantum fluctuations perturbatively in the spin-length, approaching from the classical limit $\cramped{S\rightarrow\infty}$. This can be done through a quantum linear spin-wave analysis which introduces quantum fluctuations atop the classical ground state. This allows one to examine how accidental classical ground-state degeneracies may be lifted by these quantum fluctuations. As mentioned in Sec. V.1, we discuss in Appendix. B the formalism for a quantum linear spin-wave analysis for various regimes of the phase angle $\xi$. State selection via ObD due to quantum fluctuations at zero temperature for a particular $\xi$ is determined by the zero-point energy of the linear spin waves about an accidentally degenerate ground state parametrized by angle $\phi$, $\epsilon_{\rm{Q}}(\phi)=(1/2)\sum_{\bm{q}}\omega_{\bm{q}}(\phi)$. As mentioned previously in Sec. V.1, $\omega_{\bm{q}}(\phi)$ implicitly depends on $\xi$. The accidental ground state for which the zero-point energy is minimized is selected by quantum fluctuations, resulting in quantum ObD. We find that in Regime-I, II, IV, and V, the zero-point energy distinguishes between different accidentally degenerate classical ground states and selects an ordering pattern. Interestingly, in each of the above four regimes, the zero-point energy is found to be minimized for the very same set of states chosen by the classical thermal ObD mechanism at low temperatures. Thus, ObD from quantum fluctuations at zero temperature predicts the same phase diagram as was found from the classical low-temperature expansion.
@@ VI.2 Order by disorder from combined quantum and thermal fluctuations at $T>0$
We have found that quantum ObD at zero temperature, and classical ObD from thermal fluctuations at low temperatures select the same long-range magnetic orders. However, we have not yet investigated ObD from combined quantum and thermal fluctuations at small non-zero temperatures. This regime would appear when $\cramped{T\sim O(\omega_{\bm{q}})}$; between the low-temperature quantum limit ($\cramped{T\ll\omega_{\bm{q}}}$) and the classical spin-wave limit ($\cramped{\omega_{\bm{q}}\ll T\ll T_{N}}$) as discussed in Section V.1  [^6] In typical spin systems where the spin-length lies between $S=1/2$ and $S=7/2$, not all of these regimes will be practically achievable.. ObD state selection at zero temperature and non-zero temperatures at $O(1/S)$ do not necessarily need to be the same, as was found (e.g.) for the nearest-neighbor Heisenberg fcc antiferromagnet Schick et al. (2020). Therefore, it is important to explore ObD state selection at $\cramped{T>0}$ including both quantum and thermal fluctuations. For this purpose, we focus on the free energy of the quantum linear spin waves, which is given by Kardar (2007)
@@
with $\omega_{\bm{q}}(\phi)$ being the known spin-wave spectrum about an accidentally degenerate ground state characterized by $\phi$ for a given $\xi$. The first and second terms in Eq. (9) correspond to the zero-point energy, $\epsilon_{\rm Q}(\phi)$, and the non-zero temperature contributions of the quantum spin waves to the free energy, respectively. ObD selects the states for which the free energy [Eq. (9)] is minimized. We numerically compute the free energy as a function of $\phi$ for any $\xi$ in Regime-I and Regime-IV, and find that it is minimized for the same states as those selected by the zero-point energy alone, described in Sec. VI.1. That is, the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ for any coupling parameter in Regime-I, and the ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ for any coupling parameter in Regime-IV get selected by ObD at non-zero temperatures. By using the Klein duality, Stripe-$\parallel$ and Stripe-$\perp$ are selected by ObD in Regime-II and V, respectively.
@@
In summary, quantum ObD at zero temperature, quantum-thermal ObD at non-zero temperature, and classical thermal ObD all select the same states. In other words, as has been found in many systems exhibiting ObD Henley (1989); Chaloupka et al. (2010); Price and Perkins (2012); Savary et al. (2012); McClarty et al. (2014), but is not guaranteed Schick et al. (2020), quantum and thermal fluctuations do not compete in their respective selection of ground states via ObD.
@@
Figure 4: Ground state energy per spin, $\varepsilon_{0}\equiv E_{0}/N$ where $E_{0}$ is the ground state energy eigenvalue, and the second derivative of $\varepsilon_{0}$ with respect to the phase angle $\xi$ for $L=4$. The spikes in the second derivative provide the quantum phase boundaries, exhibiting excellent agreement with the classical phase boundaries (grey dashed lines). Regime-I to VI are labelled in accordance with Fig. 2.
@@
Figure 5: (a) Results obtained from exact diagonalization of the quantum Heisenberg-compass model [Eq. (1)] on the $4\times 4$ square lattice with periodic boundary conditions. Several static structure factors [Eq. (10)] computed over the full range of $\xi$ at the four high symmetry points of the Brillouin zone, $\mathbf{\Gamma}=(0,0),\mathbf{X}=(\pi,0),\mathbf{Y}=(0,\pi)$, and $\mathbf{M}=(\pi,\pi)$. Well-defined crossovers in the dominant structure factor are in excellent agreement with the classical phase boundaries (grey dashed lines). Regime-I to VI are labelled in accordance with Fig. 2. (b) Ratio of the ground state overlap with the Néel coherent state along $\phi$, $\Omega_{\rm AFM}(\phi)$, to the maximum of the overlap over the full range of $\phi$ with $\xi=0.25\pi$ for several system sizes in Regime-I. (c) Ratio of the ground state overlap with the ferromagnetic coherent state along $\phi$, $\Omega_{\rm{FM}}(\phi)$, to the maximum of the overlap over the full range of $\phi$ with $\xi=1.25\pi$ for several system sizes in Regime-IV. The range of $\phi$ is restricted to be only $[0,\pi/2)$ due to the $C_{4}$ symmetry of the model.
@@ VI.3 Numerical exact diagonalization
Given the material examples of interest have spin $S=1/2$, the semi-classical results derived at large-$S$ must be corroborated by direct calculations in the small-$S$ limit. To this end, we investigate the quantum spin-$1/2$ Heisenberg-compass model [Eq. (1)] using exact diagonalization Sandvik (2010) on small clusters with periodic boundary conditions. For a system of $N=L^{2}$ spins, the dimension of the Hilbert space is given by $2^{N}$, which grows rapidly with system size and limits the size of cluster that can be realistically considered. To consider clusters that are as large as possible, we exploit the discrete translation symmetry of the model which divides the Hilbert space into $L^{2}$ momentum sectors which block diagonalize the Hamiltonian (see Ref. Sandvik (2010) for details). These blocks can be diagonalized individually using sparse diagonalization methods Wu and Simon (2000), such as the Lanczos algorithm, which can efficiently extract the low-lying energy eigenvalues and eigenvectors. Using this approach we are able to consider system sizes up to $L=5$ ($N=25$) with reasonable computational effort. Since the ground state orderings that we expect in Regime-I, II, V, and VI have two sublattice magnetic unit cells, these are only compatible with even values of $L$, limiting a global phase diagram to only $L=4$. However, the ferromagnetic orderings are compatible with the $L=5$ case and thus Regimes III and IV can be studied at this larger size. To capture the full phase diagram, we primarily present the results obtained using exact diagonalization performed on a square lattice of size $L=4$, which is compatible with the magnetic ordering in all parameter regimes, unless otherwise specified.
@@
We first determine the ground state and its energy, $E_{0}$, for a range of $\xi$ values across the full parameter space for $L=4$. In Fig. 4, we illustrate how the ground state energy per spin, $\varepsilon_{0}={E_{0}}/{N}$ varies with $\xi$. Near a quantum phase transition, we expect the ground state energy to change sharply with $\xi$ Sachdev (2011) and thus to identify changes in the ground state phase, we consider not just $\varepsilon_{0}$ itself, but also its derivatives with respect to $\xi$. Explicitly, to clearly identify the points at which $\varepsilon_{0}$ is changing quickly, we numerically computed the second derivative, $-{\partial^{2}\varepsilon_{0}}/{\partial\xi^{2}}$  [^7] The second derivative is computed numerically with Python’s NumPy gradient function using energy eigenvalues extracted from the numerical exact diagonalization.. This quantity is expected to show sharp peaks near any phase transition; for example, for a level crossing (first order transition) we would expect a discontinuity in the first derivative of $\varepsilon_{0}$ and thus a $\delta$-function in the second derivative. Consequently, the peaks in $-{\partial^{2}\varepsilon_{0}}/{\partial\xi^{2}}$ [shown in Fig. 4] provide good indicators for the locations of the boundaries between different phases. As depicted in Fig. 4, $\varepsilon_{0}$ exhibits pronounced kinks near $\xi=\pi/2,\pi$, and HF*. Looking at the second derivative, we find that there are pronounced peaks at $\xi=\pi/2,\pi$ and HF*, suggesting transitions at those points, and less pronounced peaks at $\xi=0,$ HAF*, and $23\pi/$ 2, suggesting weaker transitions Chaloupka et al. (2010). The location of these peaks is in excellent agreement with the phase boundaries obtained from the classical or semi-classical methods, and what one expects based on the constraints from the Klein duality. Therefore, given the system sizes for which we are able to perform exact diagonalization, the same number of phases are observed in the quantum spin-1/2 version of the Heisenberg-compass model as seen in its classical counterpart.
@@
$\mathcal{S}_{\mu\mu}(\bm{q})=\frac{1}{N}\sum_{i,j}e^{-i\bm{q}\cdot(\bm{r}_{i}-\bm{r}_{j})}\langle S_{i}^{\mu}S_{j}^{\mu}\rangle,$ (10)
@@
The nature of the phases in the spin-1/2 model near the compass points ($\xi=\pi/2$ and $23\pi/$ 2) deserves some discussion. An earlier exact diagonalization study Dorier et al. (2005) has found a set of $2^{(L+1)}$ low lying states that collapse into degenerate ground states as the system size is increased. A quantum Monte Carlo study Wenzel and Janke (2008) reveals that the spin-1/2 case also exhibits a directional or nematic ordering transition at finite temperature, similarly to the classical result. We are not aware of any detailed study of the robustness of the directional-ordered phase at non-zero temperature against weak symmetric perturbations such as the Heisenberg interaction considered here. How these perturbations affect the phase in the compass limit – and in particular its stability – is a question we leave to future work.
@@
where $\ket{\uparrow}$, $\ket{\downarrow}$ are $\bm{\hat{z}}$ quantized spin-1/2 states and $\bm{r}$ labels the sites of the lattice. The overlap of this Néel state with |$\ket{\Omega}$, that we refer to as $\Omega(\phi)_{\rm AFM}\equiv|\braket{\Omega}{\phi}_{\rm AFM}|^{2}$, for $\xi=\pi/4$ is presented in Fig. 5(b), exhibiting maximal overlap with the Néel states corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ (i.e., the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$). We also find that this overlap increases as the system size increases. We note that while $\Omega(\phi)_{\rm AFM}$ does not explicitly represent the probability of finding the Néel coherent state along $\phi$ in the ground state due to the non-orthogonality of different coherent states, this quantity gives us a qualitative idea of how close the ground state is to a given product state. We find this to be true for any value of $\xi$ in Regime-I. This thus confirms the preference for ObD-selected Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ within the quantum ground state in this regime. Note the smallness of the change in the overlaps as a function of angle $\phi$ likely originates from the smallness of the ObD selection, which is several orders of magnitude smaller than the scale of $J$ or $K$.
@@
defining $\Omega(\phi)_{\rm FM}\equiv|\braket{\Omega}{\phi}_{\rm FM}|^{2}$ where |$\ket{\Omega}$ now represents the ground state wavefunction for a given $\xi$ in Regime-IV obtained from exact diagonalization. We show the results for various system sizes with $\xi=1.25\pi$ in Fig. 5(c). The overlap is found to be maximal for the in-plane angles $\phi=0,\pi/2,\pi,3\pi/2$, confirming the preference for the ObD selected ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ within the quantum ground state. While Fig. 5(c) is shown for a particular $\xi$ [i.e., $\xi=1.25\pi$], the same is true for any $\xi$ in Regime-IV. Using the Klein duality, we argue that the ObD-selected product states have dominant contributions to the quantum ground state wavefunctions in Regime-II and V, as well. Therefore, the quantum ground state phase diagram found from exact diagonalization on small systems like $4\times 4$ square lattice is qualitatively similar to the low-temperature classical phase diagram and zero temperature semi-classical phase diagram discussed previously.

### v2

@@ VI Quantum ground state phase diagram
We have described above the classical phase diagram of the Heisenberg-compass model using several complementary methods: Luttinger-Tisza analysis, low-temperature spin-wave expansion and Monte Carlo simulations. In this section, we explore the quantum ground state phase diagram of this model. To begin, we examine the role of quantum fluctuations on the classical ground states at zero temperature. We commit particular attention to the regions where the classical model displays an accidental degeneracy and quantum ObD is expected to determine the ordering pattern.
@@ VI.1 Order by disorder from quantum fluctuations at $T=0$
We start by considering the effects of quantum fluctuations perturbatively in the spin-length, approaching from the classical limit $\cramped{S\rightarrow\infty}$. This can be done through a quantum linear spin-wave analysis which introduces quantum fluctuations atop the classical ground state. This allows one to examine how accidental classical ground-state degeneracies may be lifted by these quantum fluctuations. As mentioned in Sec. V.1, we discuss in Appendix B the formalism for a quantum linear spin-wave analysis for various regimes of the phase angle $\xi$. State selection via ObD due to quantum fluctuations at zero temperature for a particular $\xi$ is determined by the zero-point energy of the linear spin waves about an accidentally degenerate ground state parametrized by angle $\phi$, $\epsilon_{\rm{Q}}(\phi)=(1/2)\sum_{\bm{q}}\omega_{\bm{q}}(\phi)$. As mentioned previously in Sec. V.1, $\omega_{\bm{q}}(\phi)$ implicitly depends on $\xi$. The accidental ground state for which the zero-point energy is minimized is selected by quantum fluctuations, resulting in quantum ObD. We find that in Regime-I, II, IV, and V, the zero-point energy distinguishes between different accidentally degenerate classical ground states and selects a long-range ordered pattern. Interestingly, in each of the above four regimes, the zero-point energy is found to be minimized for the very same set of states chosen by the classical thermal ObD mechanism at low temperatures. Thus, ObD from quantum fluctuations at zero temperature predicts the same phase diagram as was found from the classical low-temperature expansion.
@@ VI.2 Order by disorder from combined quantum and thermal fluctuations at $T>0$
We have found that quantum ObD at zero temperature, and classical ObD from thermal fluctuations at low temperatures select the same long-range magnetic orders. However, we have not yet investigated ObD from combined quantum and thermal fluctuations at small non-zero temperatures. This regime would appear when $\cramped{T\sim O(\omega_{\bm{q}})}$; between the low-temperature quantum limit ($\cramped{T\ll\omega_{\bm{q}}}$) and the classical spin-wave limit ($\cramped{\omega_{\bm{q}}\ll T\ll T_{c}}$) as discussed in Section V.1  [^6] In typical spin systems where the spin-length lies between $S=1/2$ and $S=7/2$, not all of these regimes will be practically realizable.. ObD state selection at zero temperature and non-zero temperatures at $O(1/S)$ do not necessarily need to be the same, as was found in Refs. Schick et al. (2020); Noculak et al. (2023); Hickey et al. (2024). Therefore, it is important to explore ObD state selection at $\cramped{T>0}$ including both quantum and thermal fluctuations. For this purpose, we focus on the free energy of the quantum linear spin waves, which is given by Kardar (2007)
@@
with $\omega_{\bm{q}}(\phi)$ being the known spin-wave spectrum about an accidentally degenerate ground state characterized by $\phi$ for a given $\xi$. The first and second terms in Eq. (9) correspond to the zero-point energy, $\epsilon_{\rm Q}(\phi)$, and the non-zero temperature contributions of the quantum spin waves to the free energy, respectively. ObD selects the states for which the free energy [Eq. (9)] is minimized. We numerically compute the free energy as a function of $\phi$ for any $\xi$ in Regime-I and Regime-IV, and find that it is minimized for the same states as those selected by the zero-point energy alone, described in Sec. VI.1. That is, the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ for any coupling parameter in Regime-I, and the ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ for any coupling parameter in Regime-IV get selected by ObD at non-zero temperatures. By using the Klein duality, we conclude thatStripe-$\parallel$ and Stripe-$\perp$ are selected by ObD in Regime-II and V, respectively.
@@
In summary, quantum ObD at zero temperature, quantum-thermal ObD at non-zero temperature, and classical thermal ObD all select the same states. In other words, as has been found in many systems exhibiting ObD Henley (1989); Chaloupka et al. (2010); Price and Perkins (2012); Savary et al. (2012); McClarty et al. (2014), but is not guaranteed Schick et al. (2020); Noculak et al. (2023); Hickey et al. (2024), quantum and thermal fluctuations do not compete in their respective selection of ground states via ObD.
@@
Figure 4: Ground state energy per spin, $\varepsilon_{0}\equiv E_{0}/N$ where $E_{0}$ is the ground state energy eigenvalue, and the second derivative of $\varepsilon_{0}$ with respect to the phase angle $\xi$ for $L=4$. The second derivative is computed numerically using a finite-difference formula using $\varepsilon_{0}$ obtained from exact diagonalization. The spikes in the second derivative provide the quantum phase boundaries, exhibiting excellent agreement with the classical phase boundaries (grey dashed lines). Regime-I to VI are labelled in accordance with Fig. 2.
@@
Figure 5: (a) Results obtained from exact diagonalization of the quantum Heisenberg-compass model [Eq. (1)] on the $4\times 4$ square lattice with periodic boundary conditions. Several static structure factors [Eq. (10)] computed over the full range of $\xi$ at the four high symmetry points of the Brillouin zone, $\mathbf{\Gamma}=(0,0),\mathbf{X}=(\pi,0),\mathbf{Y}=(0,\pi)$, and $\mathbf{M}=(\pi,\pi)$. Well-defined crossovers in the dominant structure factor are in excellent agreement with the classical phase boundaries (grey dashed lines). Regime-I to VI are labelled in accordance with Fig. 2. (b) Ratio of the ground state overlap with the Néel coherent state along $\phi$, $\Omega_{\rm AFM}(\phi)$, to the maximum of the overlap over the full range of $\phi$, $\Omega_{\rm max}$, with $\xi=0.25\pi$ for several system sizes in Regime-I. (c) Ratio of the ground state overlap with the ferromagnetic coherent state along $\phi$, $\Omega_{\rm{FM}}(\phi)$, to the maximum of the overlap over the full range of $\phi$, $\Omega_{\rm max}$, with $\xi=1.25\pi$ for several system sizes in Regime-IV. The range of $\phi$ is restricted to be only $[0,\pi/2)$ due to the $C_{4}$ symmetry of the model.
@@ VI.3 Numerical exact diagonalization
Given the material examples of interest have spin $S=1/2$, the semi-classical results derived at large-$S$ must be corroborated by direct calculations in the small-$S$ limit. To this end, we investigate the quantum spin-$1/2$ Heisenberg-compass model [Eq. (1)] using exact diagonalization Sandvik (2010) on small clusters with periodic boundary conditions. For a system of $N=L^{2}$ spins, the dimension of the Hilbert space is given by $2^{N}$, which grows rapidly with system size and limits the size of cluster that can be realistically considered. To consider clusters that are as large as possible, we exploit the discrete translation symmetry of the model which divides the Hilbert space into $L^{2}$ momentum sectors which block diagonalize the Hamiltonian (see Ref. Sandvik (2010) for details). These blocks can be diagonalized individually using sparse diagonalization methods Wu and Simon (2000), such as the Lanczos algorithm, which can efficiently extract the low-lying energy eigenvalues and eigenvectors. Using this approach we are in principle able to consider system sizes up to $L=5$ ($N=25$) with reasonable computational effort. However, since the ground state orderings that we expect in Regime-I, II, V, and VI have two sublattice magnetic unit cells, these are only compatible with even values of $L$, limiting a global phase diagram to only $L=4$. However, the ferromagnetic orderings are compatible with the $L=5$ case and thus Regimes III and IV can be studied at this larger size. To capture the full phase diagram, we primarily present the results obtained using exact diagonalization performed on a square lattice of size $L=4$, which is compatible with the magnetic ordering in all parameter regimes, unless otherwise specified.
@@
We first determine the ground state and its energy, $E_{0}$, for a range of $\xi$ values across the full parameter space for $L=4$. In Fig. 4, we illustrate how the ground state energy per spin, $\varepsilon_{0}={E_{0}}/{N}$ varies with $\xi$. Near a quantum phase transition, we expect the ground state energy to change sharply with $\xi$ Sachdev (2011) and thus to identify changes in the ground state phase, we consider not just $\varepsilon_{0}$ itself, but also its derivatives with respect to $\xi$. Explicitly, to clearly identify the points at which $\varepsilon_{0}$ is changing quickly, we numerically computed the second derivative, $-{\partial^{2}\varepsilon_{0}}/{\partial\xi^{2}}$ using a finite-difference formula. This quantity is expected to show sharp peaks near any phase transition; for example, for a level crossing (first order transition) we would expect a discontinuity in the first derivative of $\varepsilon_{0}$ and thus a $\delta$-function in the second derivative. Consequently, the peaks in $-{\partial^{2}\varepsilon_{0}}/{\partial\xi^{2}}$ [shown in Fig. 4] provide good indicators for the locations of the boundaries between different phases. As depicted in Fig. 4, $\varepsilon_{0}$ exhibits pronounced kinks near $\xi=\pi/2,\pi$, and HF*. Looking at the second derivative, we find that there are pronounced peaks at $\xi=\pi/2,\pi$ and HF*, suggesting transitions at those points, and less pronounced peaks at $\xi=0,$ HAF*, and $23\pi/$ 2, suggesting weaker transitions Chaloupka et al. (2010). Note that the peak in the second derivative between Regime-II and Regime-III is not exactly at HAF*, but slightly shifted to the right. However, Klein duality dictates that if HAF is a phase boundary, so is HAF*. That HAF is a phase boundary can be understood in both the quantum and classical cases through the breaking of the SU(2) spin rotation symmetry –- picking either an XY (in Regime-I) or Ising-like (in Regime-VI) ordered phase near the HAF point. Therefore, the peak in $-{\partial^{2}\varepsilon_{0}}/{\partial\xi^{2}}$ in Fig. 4 between Regime-II and Regime-III is expected to be exactly at HAF*. On the basis of the argument above, we believe the discrepancy between the peak position and HAF* is not due to any (peculiar) quantum effects but, rather, it is likely due to finite size effects, which should diminish as the system size increases. Apart from this slight discrepancy, the locations of the peaks are in excellent agreement with the phase boundaries obtained from the classical or semi-classical methods, and what one expects based on the constraints from the Klein duality. Therefore, given the system sizes for which we are able to perform exact diagonalization, the same number of phases are observed in the quantum spin-1/2 version of the Heisenberg-compass model as seen in its classical counterpart.
@@
$\mathcal{S}_{\mu\mu}(\bm{q})=\frac{1}{N}\sum_{\bm{r},\,\bm{r}^{\prime}}e^{-i\bm{q}\cdot(\bm{r}-\bm{r}^{\prime})}\langle S_{\bm{r}\phantom{{}^{\prime}}}^{\mu}S_{\bm{r}^{\prime}}^{\mu}\rangle,$ (10)
@@
The nature of the phases in the spin-1/2 model near the compass points ($\xi=\pi/2$ and $23\pi/$ 2) deserves some discussion. At the compass points, an exact diagonalization study Dorier et al. (2005) has found a set of $2^{(L+1)}$ low lying states that collapse into degenerate ground states as the system size is increased. A quantum Monte Carlo study Wenzel and Janke (2008) reveals that the spin-1/2 case also exhibits a directional or nematic ordering transition at finite temperature, similarly to the classical result. We are not aware of any detailed study of the robustness of the directional-ordered phase at non-zero temperature against weak symmetric perturbations such as the Heisenberg interaction considered here. How these perturbations affect the directional-ordered phase in the vicinity of the pure compass limit – and in particular its stability – is a question we leave for future work.
@@
where $\ket{\uparrow}$, $\ket{\downarrow}$ are $\bm{\hat{z}}$ quantized spin-1/2 states and $\bm{r}$ labels the sites of the lattice. The overlap of this Néel state with |$\ket{\Omega}$, that we refer to as $\Omega_{\rm AFM}(\phi)\equiv|\braket{\Omega}{\phi}_{\rm AFM}|^{2}$, for $\xi=\pi/4$ is presented in Fig. 5(b), exhibiting maximal overlap with the Néel states corresponding to $\phi=0,\pi/2,\pi,3\pi/2$ (i.e., the Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$). We also find that the anisotropy of this overlap increases as the system size increases. We note that while $\Omega_{\rm AFM}(\phi)$ does not explicitly represent the probability of finding the Néel coherent state along $\phi$ in the ground state due to the non-orthogonality of different coherent states, this quantity gives us a qualitative idea of how close the ground state is to a given product state. We find this to be true for any value of $\xi$ in Regime-I. This thus confirms the preference for ObD-selected Néel states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ within the quantum ground state in this regime. Note the smallness of the change in the overlaps as a function of angle $\phi$ likely originates from the weakness of the ObD selection (discussed in Sec. VI.1), which is several orders of magnitude smaller than the scale of $J$ or $K$.
@@
defining $\Omega_{\rm FM}(\phi)\equiv|\braket{\Omega}{\phi}_{\rm FM}|^{2}$ where |$\ket{\Omega}$ now represents the ground state wavefunction for a given $\xi$ in Regime-IV obtained from exact diagonalization. We show the results for various system sizes with $\xi=1.25\pi$ in Fig. 5(c). The overlap is found to be maximal for the in-plane angles $\phi=0,\pi/2,\pi,3\pi/2$, confirming the preference for the ObD selected ferromagnetic states along ±$\pm\bm{\hat{x}},\pm\bm{\hat{y}}$ within the quantum ground state. While Fig. 5(c) is shown for a particular $\xi$ [i.e., $\xi=1.25\pi$], the same was found to be true for any $\xi$ in Regime-IV. Using the Klein duality, we argue that the ObD-selected product states have dominant contributions to the quantum ground state wavefunctions in Regime-II and V, as well. Therefore, the quantum ground state phase diagram found from exact diagonalization on small systems like $4\times 4$ square lattice is qualitatively similar to the low-temperature classical phase diagram and zero temperature semi-classical phase diagram discussed previously.